OpenStudy (anonymous):

The probability that a dessert sold at a certain cafe contains chocolate is 73%. The probability that a dessert containing chocolate also contains nuts is 25%. Find the probability that a dessert chosen at random contains nuts given that it contains chocolate. Round to the nearest tenth of a percent.

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Is my aswer 48 ?

OpenStudy (anonymous):

Whoops wrong question. This is the one i meant to post !

OpenStudy (anonymous):

An airline has 81% of its flights depart on schedule. It has 69% of its flights depart and arrive on schedule. Find the probability that a flight that departs on schedule also arrives on schedule.

jimthompson5910 (jim_thompson5910):

P(Contains Nuts| Chocolate) = P(Chocolate and nuts)/P(Chocolate) P(Contains Nuts| Chocolate) = 0.25/0.73 P(Contains Nuts| Chocolate) = 0.34246575342466

OpenStudy (anonymous):

I posted the wrong question . Its above of what you said ^ ?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

P(Arrive On Schedule|Depart On Schedule) = P(Depart and Arrive On Schedule)/P(Depart On Schedule) P(Arrive On Schedule|Depart On Schedule) = 0.69/0.81 P(Arrive On Schedule|Depart On Schedule) = 0.85185185185186

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Find the values of the 30th and 90th percentiles of the data. 18, 9, 7, 5, 11, 7, 17, 20, 19, 2, 17, 12, 5, 1, 13, 12, 11, 15, 16, 20

OpenStudy (anonymous):

I got 30th percentile = 9 90th percentile = 20 but that answer isnt on my answer choice so what could it be ?

jimthompson5910 (jim_thompson5910):

what do you get when you sort the data

OpenStudy (anonymous):

Here's my asnwer choice 30th percentile = 7 90th percentile = 19 30th percentile = 9 90th percentile = 19 30th percentile = 11 90th percentile = 19 30th percentile = 7 90th percentile = 20

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

You there ?

jimthompson5910 (jim_thompson5910):

what do you get when you sort the data

OpenStudy (anonymous):

What do you mean ?

jimthompson5910 (jim_thompson5910):

when you arrange the list from smallest number to largest number, what do you get?

OpenStudy (anonymous):

Huh ? I just posted the question. After doing the problem i got I got 30th percentile = 9 90th percentile = 20 But that answer isnt on my answer problem. So i posted you the answer choice to see what one is really my answer,

jimthompson5910 (jim_thompson5910):

example the list 3, 1, 2 is not in order but sorting them gives you 1, 2, 3

OpenStudy (anonymous):

exactly

jimthompson5910 (jim_thompson5910):

so sort the given list in your problem to get _____

OpenStudy (anonymous):

After i sort them then whats my answer ?

jimthompson5910 (jim_thompson5910):

I'll show you how to find it once you've sorted it

OpenStudy (anonymous):

I have the correct answer just that correct answer isn't on my answer choice.

OpenStudy (anonymous):

is my answer 30th percentile = 9 90th percentile = 19 or 30th percentile = 7 90th percentile = 20

jimthompson5910 (jim_thompson5910):

the sorted list is 1, 2, 5, 5, 7, 7, 9, 11, 11, 12, 12, 13, 15, 16, 17, 17, 18, 19, 20, 20

jimthompson5910 (jim_thompson5910):

there are 20 values in this list so the 30th percentile is at location number 0.3*20 = 6

jimthompson5910 (jim_thompson5910):

So the the 30th percentile is at location number 6

OpenStudy (anonymous):

There is no 6

jimthompson5910 (jim_thompson5910):

slot number 6

OpenStudy (anonymous):

Here's my asnwer choice 30th percentile = 7 90th percentile = 19 30th percentile = 9 90th percentile = 19 30th percentile = 11 90th percentile = 19 30th percentile = 7 90th percentile = 20

jimthompson5910 (jim_thompson5910):

which number is in slot 6

jimthompson5910 (jim_thompson5910):

i know

OpenStudy (anonymous):

7

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

the number 7 is in slot 6 (6th spot from the left)

jimthompson5910 (jim_thompson5910):

so 7 is the 30th percentile

OpenStudy (anonymous):

whats 90th ? 19 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

30th percentile = 7 90th percentile = 19

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

correct ?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you. I have a few more,.

jimthompson5910 (jim_thompson5910):

ok just a few though

OpenStudy (anonymous):

Find the mean, median, and mode of the set of data 10, 11, 4, 7, 12, 11, 16, 6, 9, 15

jimthompson5910 (jim_thompson5910):

mean = add up the numbers, then divide by 10 median = middle most number (after you sort the numbers) mode = most occurring number

OpenStudy (anonymous):

Thats it ?

jimthompson5910 (jim_thompson5910):

tell me what you get when you follow those steps

OpenStudy (anonymous):

Oh the mean is were i add them all

jimthompson5910 (jim_thompson5910):

and divide by 10

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

10.1

jimthompson5910 (jim_thompson5910):

that's the mean, good

OpenStudy (anonymous):

okay . . .

OpenStudy (anonymous):

12 and 11 are in the middle

jimthompson5910 (jim_thompson5910):

sort the data to get 4, 6, 7, 9, 10, 11, 11, 12, 15, 16

OpenStudy (anonymous):

kidding 11

jimthompson5910 (jim_thompson5910):

the middle most numbers are: 10, 11 average them to find the median

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

add subtract divide what ?

jimthompson5910 (jim_thompson5910):

add them up, divide by 2

OpenStudy (anonymous):

so my median is 10.5

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

the mode is the value that shows up the most

OpenStudy (anonymous):

11 ?

jimthompson5910 (jim_thompson5910):

which number shows up the most

jimthompson5910 (jim_thompson5910):

yep, 11

jimthompson5910 (jim_thompson5910):

it shows up 2 times (everything else shows up 1 time)

OpenStudy (anonymous):

Lucas recorded his lunch expenditure each day for one week in the table below.

OpenStudy (anonymous):

Find the mean, standard deviation, and variance of Lucas’ lunch expenditures. Round to the nearest thousandth.

jimthompson5910 (jim_thompson5910):

mean = add up the numbers and divide by 7

OpenStudy (anonymous):

5.1

jimthompson5910 (jim_thompson5910):

yep that's the mean

jimthompson5910 (jim_thompson5910):

now subtract the mean from each value

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

4.85-5.1 = -0.25 5.10-5.1 = 0 5.50-5.1 = 0.4 4.75-5.1 = -0.35 4.50-5.1 = -0.6 5.00-5.1 = -0.1 6.00-5.1 = 0.9

jimthompson5910 (jim_thompson5910):

So the differences are -0.25, 0, 0.4, -0.35, -0.6, -0.1, 0.9

OpenStudy (anonymous):

is that it for my problem or answer ?

jimthompson5910 (jim_thompson5910):

Square each difference to get (-0.25)^2 = 0.0625 (0)^2 = 0 (0.4)^2 = 0.16 (-0.35)^2 = 0.1225 (-0.6)^2 = 0.36 (-0.1)^2 = 0.01 (0.9)^2 = 0.81

jimthompson5910 (jim_thompson5910):

So the list of squares are 0.0625, 0, 0.16, 0.1225, 0.36, 0.01, 0.81

jimthompson5910 (jim_thompson5910):

Add them up to get 0.0625+0+0.16+0.1225+0.36+0.01+0.81 = 1.525 Divide by 6 to get 1.525/6 = 0.25416666666667 and finally take the square root of that to get sqrt(0.25416666666667) = 0.5041494487418

OpenStudy (anonymous):

Is that it ? for my problem ?

jimthompson5910 (jim_thompson5910):

yeah I'm showing you how to find the variance and standard deviation

OpenStudy (anonymous):

whats my varicance

jimthompson5910 (jim_thompson5910):

that would be 0.25416666666667 and I showed how you find it above

OpenStudy (anonymous):

Identify the type of sampling method used. A trucking company places their office phone number on the back of all of their vehicles to receive comments on how well their employees are driving.

OpenStudy (anonymous):

i have 4 more thank you sooo much

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what are your choices for this one

OpenStudy (anonymous):

there are no choices

jimthompson5910 (jim_thompson5910):

hmm not sure